# Bruce W. Brewer    Haitao Wang

### v1

@@ 1 Introduction
The problem was studied previously. A straightforward solution is to first construct $G(S)$ explicitly and then run a breadth-first-search algorithm, which takes $\Omega(n^{2})$ time since $G(S)$ may have $\Omega(n^{2})$ edges in the worst case. Kaplan, Katz, Saban, and Sharir [15] proposed the first subquadratic time algorithm, and their algorithm runs in $O(n\log^{4}n)$ time. Later Klost [18] gave an $O(n\log^{2}n)$ time improved solution. On the other hand, it is known that the problem has an $\Omega(n\log n)$ time lower bound under the algebraic decision tree model even if all disks have the same radius [6].
@@ 3.3 Algorithm correctness
For any site $v\in S\setminus S_{\leq i-1}$, $v$ is in $S_{i}$ if and only if $d_{S_{i-1}}(v)\leq r_{v}$.
@@
We next prove that $S_{i}^{\prime}\supseteq S_{i}$. Consider a site $v\in S_{i}$. Our goal is to show that $v$ is also in $S_{i}^{\prime}$. We will prove in Lemma 3.1 that there exists a vertex $u\in S_{i-1}$ and a $u$-$v$ path $\pi$ in $\mathcal{DT}(S)$ such that all internal vertices of $\pi$ are in $S_{i}$ (note that internal vertices refer to vertices of $\pi$ excluding $u$ and $v$). Let the vertices of $\pi$ be $u=w_{0},w_{1},\ldots,w_{k}=v$. We argue that all $w_{j}$ with $1\leq j\leq k$ are added to $S_{i}^{\prime}$. To this end, because $w_{j}\in S_{i}$ for all $1\leq j\leq k$, it is sufficient to argue that $w_{j}$ is added to $Q$ because when $w_{j}$ is popped from $Q$, it will pass the check on Line 1 by Section 3.3 and then be added to $S_{i}^{\prime}$. Indeed, $w_{1}$ is added to $Q$ on Line 1 because $w_{1}\in N_{\mathcal{DT}}(u)$. Since $w_{2}\in N_{\mathcal{DT}}(w_{1})$, $w_{2}$ will be added to $Q$ on Line 1 no later than right after $w_{1}$ is added to $S^{\prime}_{i}$ on Line 1. Following the same argument, $w_{3},w_{4},\ldots,w_{k}=v$ will all be added to $Q$ and thus added to $S_{i}^{\prime}$ as well. This proves that $v\in S_{i}^{\prime}$. Therefore, $S_{i}^{\prime}\supseteq S_{i}$.
@@
Let $u=\beta_{v}(S_{i-1})$, i.e., $u$ is a nearest neighbor of $v$ in $S_{i-1}$. Consider the line segment $\overline{uv}$ and the sites whose Voronoi regions in $\mathcal{VD}(S)$ intersect $\overline{uv}$. Let $w_{0},w_{1},\ldots,w_{k}$ be these sites in the order that their Voronoi regions intersect $\overline{uv}$ with $w_{0}=u$ and $w_{k}=v$. Note that repetitions are possible, so it could be that $w_{j}=w_{j^{\prime}}$ for $j\neq j^{\prime}$. See Figure 1 for an example. For ease of discussion, we assume that $\overline{uv}$ does not contain any Voronoi vertices of $\mathcal{VD}(S)$ (otherwise, we could replace $v$ by a point arbitrarily close to $v$ and then follow the same argument). This implies that, for every $10\leq j\leq k-$ 1, $\overline{uv}$ crosses the Voronoi edge between $w_{j}$ and $w_{j+1}$. This further implies that $w_{j}$ and $w_{j+1}$ are adjacent in $\mathcal{DT}(S)$, so $\pi=w_{0},w_{1},\ldots,w_{k}$ is a $u$-$v$ path in $\mathcal{DT}(S)$.
Next, we need to show that all internal vertices $w_{j}$ ($11\leq j\leq k-$ 1) of $\pi$ are in $S_{i}$. We will first prove that there is an edge $(u,w_{j})$ in $G(S)$ connecting $u$ and $w_{j}$, meaning that $w_{j}\in S_{i-2}\cup S_{i-1}\cup S_{i}$ since $u\in S_{i-1}$. Then, we will show that $w_{j}\notin S_{i-2}\cup S_{i-1}$, so $w_{j}\in S_{i}$ must hold. This will complete the proof of the lemma.
@@
Consider any $w_{j}$ with $11\leq j\leq k-$ 1. We now argue that $G(S)$ has an edge $(u,w_{j})$. Recall that $\overline{uv}$ intersects $\text{Vor}(w_{j})$. Let $p$ be a point of $\overline{uv}\cap\text{Vor}(w_{j})$; see Figure 2. We can derive the following:
@@
where (1) is due to the triangle inequality, (2) is due to $p\in\text{Vor}(w_{j})$ (therefore, $|w_{j}p|-r_{w_{j}}=d_{w_{j}}(p)\leq d_{v}(p)=|vp|-r_{v}$), (3) is due to $p\in\overline{uv}$, and (4) is due to Section 3.3 and the fact that $G(S)$ has an edge $(u,v)$. Consequently, by Section 3.3, $G(S)$ has an edge $(u,w_{j})$.
@@
To prove that $w_{j}\notin S_{i-2}\cup S_{i-1}$, $11\leq j\leq k-$ 1, as before, we pick a point $p\in\overline{uv}\cap\text{Vor}(w_{j})$ and derive the following:
@@
where (1) is due to the triangle inequality and our general position assumption (because $p\in\overline{uv}$, equality occurs only if $u$, $v$, and $w_{j}$ are collinear, violating our general position assumption), (2) is due to $p\in\text{Vor}(w_{j})$ (therefore, $|w_{j}p|-r_{w_{j}}=d_{w_{j}}(p)\leq d_{u}(p)=|up|-r_{u}$), (3) is due to $p\in\overline{uv}$, and (4) is due to Section 3.3 and the fact that $G(S)$ has an edge $(u,v)$, as proven above.
Notice that $w_{j}\notin S_{i-1}$ since otherwise $d_{w_{j}}(v)<d_{u}(v)$ proved above would contradict with $u=\beta_{v}(S_{i-1})$. Also, since $G(S)$ has an edge $(w_{j},v)$ and $v\in S_{i}$, $w_{j}$ cannot be in $S_{i-2}$ (since otherwise $v$ would be in $S_{\leq i-1}$, contradicting with that $v\in S_{i}$). This proves that $w_{j}\notin S_{i-2}\cup S_{i-1}$.

### v2

@@ 1 Introduction
The problem was studied previously. A straightforward solution is to first construct $G(S)$ explicitly and then run a breadth-first-search (BFS) algorithm, which takes $\Omega(n^{2})$ time since $G(S)$ may have $\Omega(n^{2})$ edges in the worst case. As indicated in [16, 15], the problem can be solved by BFS with a dynamic nearest neighbor data structure for the weighted Euclidean distance; with the data structure in [19], a randomized algorithm of $O(n\log^{4}n)$ time can be achieved. Klost [18] gave a (deterministic) algorithm of $O(n\log^{2}n)$ time. On the other hand, it is known that the problem has an $\Omega(n\log n)$ time lower bound under the algebraic decision tree model even if all disks have the same radius [6].
@@ 3 The algorithm
We assume that no disk fully contains another disk in the input. With this assumption, every site of $S$ defines a non-empty Voronoi region in $\mathcal{VD}(S)$ [23]. We will later explain that the algorithm can be slightly modified to accommodate the situation in which the assumption does not hold.
@@ 3.3 Algorithm correctness
For any site $\in S\setminus S_{\leq i-1}, vv$ is in $S_{i}$ if and only if $d_{S_{i-1}}(v)\leq r_{v}$.
@@
We next prove that $S_{i}^{\prime}\supseteq S_{i}$. Consider a site $v\in S_{i}$. Our goal is to show that $v$ is also in $S_{i}^{\prime}$. We will prove in Lemma˜3.1 that there exists a vertex $u\in S_{i-1}$ and a $u$-$v$ path $\pi$ in $\mathcal{DT}(S)$ such that all internal vertices of $\pi$ are in $S_{i}$ (note that internal vertices refer to vertices of $\pi$ excluding $u$ and $v$). Let the vertices of $\pi$ be $u=w_{0},w_{1},\ldots,w_{k}=v$. We argue that all $w_{j}$ with $1\leq j\leq k$ are added to $S_{i}^{\prime}$. To this end, because $w_{j}\in S_{i}$ for all $1\leq j\leq k$, it is sufficient to argue that $w_{j}$ is added to $Q$ because when $w_{j}$ is popped from $Q$, it will pass the check on Line 1 by Section˜3.3 and then be added to $S_{i}^{\prime}$. Indeed, $w_{1}$ is added to $Q$ on Line 1 because $w_{1}\in N_{\mathcal{DT}}(u)$. Since $w_{2}\in N_{\mathcal{DT}}(w_{1})$, $w_{2}$ will be added to $Q$ on Line 1 no later than right after $w_{1}$ is added to $S^{\prime}_{i}$ on Line 1. Following the same argument, $w_{3},w_{4},\ldots,w_{k}=v$ will all be added to $Q$ and thus added to $S_{i}^{\prime}$ as well. This proves that $v\in S_{i}^{\prime}$. Therefore, $S_{i}^{\prime}\supseteq S_{i}$.
@@
Let $u=\beta_{v}(S_{i-1})$, i.e., $u$ is a nearest neighbor of $v$ in $S_{i-1}$. Consider the line segment $\overline{uv}$ and the sites whose Voronoi regions in $\mathcal{VD}(S)$ intersect $\overline{uv}$. Let $w_{0},w_{1},\ldots,w_{k}$ be these sites in the order that their Voronoi regions intersect $\overline{uv}$ with $w_{0}=u$ and $w_{k}=v$. Note that repetitions are possible, so it could be that $w_{j}=w_{j^{\prime}}$ for $j\neq j^{\prime}$. See Figure˜1 for an example. For ease of discussion, we assume that $\overline{uv}$ does not contain any Voronoi vertices of $\mathcal{VD}(S)$ (otherwise, we could replace $v$ by a point arbitrarily close to $v$ and then follow the same argument). This implies that, for every $0\leq j\leq k-1$, $\overline{uv}$ crosses the Voronoi edge between $w_{j}$ and $w_{j+1}$. This further implies that $w_{j}$ and $w_{j+1}$ are adjacent in $\mathcal{DT}(S)$, so $\pi=w_{0},w_{1},\ldots,w_{k}$ is a $u$-$v$ path in $\mathcal{DT}(S)$.
@@
Next, we need to show that all internal vertices $w_{j}$ ($1\leq j\leq k-1$) of $\pi$ are in $S_{i}$. We will first prove that there is an edge $(u,w_{j})$ in $G(S)$ connecting $u$ and $w_{j}$, meaning that $w_{j}\in S_{i-2}\cup S_{i-1}\cup S_{i}$ since $u\in S_{i-1}$. Then, we will show that $w_{j}\notin S_{i-2}\cup S_{i-1}$, so $w_{j}\in S_{i}$ must hold. This will complete the proof of the lemma.
@@
Consider any $w_{j}$ with $1\leq j\leq k-1$. We now argue that $G(S)$ has an edge $(u,w_{j})$. Recall that $\overline{uv}$ intersects $\text{Vor}(w_{j})$. Let $p$ be a point of $\overline{uv}\cap\text{Vor}(w_{j})$; see Figure 2. We can derive the following:
@@
where (1) is due to the triangle inequality, (2) is due to $p\in\text{Vor}(w_{j})$ (therefore, $|w_{j}p|-r_{w_{j}}=d_{w_{j}}(p)\leq d_{v}(p)=|vp|-r_{v}$), (3) is due to $p\in\overline{uv}$, and (4) is due to Section˜3.3 and the fact that $G(S)$ has an edge $(u,v)$. Consequently, by Section˜3.3, $G(S)$ has an edge $(u,w_{j})$.
@@
To prove that $w_{j}\notin S_{i-2}\cup S_{i-1}$, $1\leq j\leq k-1$, as before, we pick a point $p\in\overline{uv}\cap\text{Vor}(w_{j})$ and derive the following:
@@
where (1) is due to the triangle inequality and our general position assumption (because $p\in\overline{uv}$, equality occurs only if $u$, $v$, and $w_{j}$ are collinear, violating our general position assumption), (2) is due to $p\in\text{Vor}(w_{j})$ (therefore, $|w_{j}p|-r_{w_{j}}=d_{w_{j}}(p)\leq d_{u}(p)=|up|-r_{u}$), (3) is due to $p\in\overline{uv}$, and (4) is due to Section˜3.3 and the fact that $G(S)$ has an edge $(u,v)$, as proven above.
We find that $w_{j}\notin S_{i-1}$ since otherwise $d_{w_{j}}(v)<d_{u}(v)$ proved above would contradict with $u=\beta_{v}(S_{i-1})$. Also, since $G(S)$ has an edge $(w_{j},v)$ and $v\in S_{i}$, $w_{j}$ cannot be in $S_{i-2}$ (since otherwise $v$ would be in $S_{\leq i-1}$, contradicting with that $v\in S_{i}$). This proves that $w_{j}\notin S_{i-2}\cup S_{i-1}$.
The containment case.
The above discusses the case where no disk fully contains another disk in the input. Suppose that a disk is contained in another disk. Then, the Voronoi region of the center of the smaller disk in $\mathcal{VD}(S)$ is empty [23], and our algorithm will miss the disk. To fix the issue, we can slightly revise our algorithm as follows. Let $v$ be a site of $S$ whose Voronoi region in $\mathcal{VD}(S)$ is empty. Let $u$ be the site of $S$ whose Voronoi region contains $v$ (note that the disk of $v$ must be contained in the disk of $u$). Then, we add $v$ to $N_{\mathcal{DT}}(u)$ and also add $u$ to $N_{\mathcal{DT}}(v)$. After doing this for all such $v$, we apply exactly the same algorithm as before. One can verify that the algorithm works even if the source site $s$ defines an empty Voronoi region in $\mathcal{VD}(S)$. The total time of the algorithm is still $O(n\log n)$.